using logarithma , find 4 1/3 to the nearest tenth
Answer by stanbon(75887)

using logarithms, find 4 1/3 to the nearest tenth
----------
Let x = 4 1/3
Then x = 13/3
log(x) = log(13)-log(3)
log(x) = 0.6368
x = 10^(0.6268)
x = 4.3
==============
Cheers,
Stan H.
================